\begin{thebibliography}{99}
%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
%\bibitem{HJ91} R. A. Horn and C. R. Johnson, Topics in Matrix Analysis, Cambridge University Press, Cambridge, 1991.
%\bibitem{BP1994} A. Berman and R. J. Plemmons, Nonnegative Matrices in the Mathematical Sciences, Classics
%in Appl. Math. 9, SIAM, Philadelphia, 1994
%\bibitem{BM2003} A. Berman and N. Shaked-Monderer, Completely Positive Matrices, World Scientific, New Jersy, 2003
\bibitem{CCL2015} A. Chang, J. Cooper and W. Li, Analytic connectivity of $k$-uniform hypergraphs, arXiv:1507.02763 (2015)
\bibitem{CQZ2013} K.C. Chang, L. Qi and T. Zhang, A survey on the spectral theory of
nonnegative tensors, Numerical Linear Algebra with Applications, 20 (2013) 891-912
%\bibitem{CQW} M. Che, L. Qi and Y. Wei, Positive definite tensors to nonlinear complemtarity problems, arXiv:1501.02546, 2015
%\bibitem{CLQ2015} H. Chen, G. Li and L. Qi, Further results on Cauchy tensors and Hankel tensors, arXiv:1501.06726v2, 2015
%\bibitem{CQ2015} H. Chen and L. Qi, Positive definiteness and semi-definiteness of even order symmetric Cauchy tensors, Journal of Industrial and Management Optimization, 11(4)(2015) 1263-1274
%\bibitem{CQi2015} Z. Chen and L. Qi, Circulant tensors with applications to spectral hypergraph theory and stochastic process, April 2014   arXiv:1312.2752
\bibitem{CD2012} J. Cooper and A. Dutle, Spectra of uniform hypergraphs, Linear Algebra and Its Applications 436 (2012) 3268-3292
%\bibitem{CZPA}
%A. Cichocki, R. Zdunek, A.H. Phan and S. Amari, {\it Nonnegative
%Matrix and Tensor Factorizations: Applications to Exploratory
%Multiway Data Analysis and Blind Source Separation}, Wiley, New
%York, 2009.
%\bibitem{CLQ2015a} Haibin Chen, Guoyin Li and Liqun Qi, Sum-of-squares tensors and their sum-of-squares rank, 2015 arXiv:submit/1230607
%\bibitem{Dickinson2010} P.I.C. Dickinson, An improved characterization of the interior of the completely positive cone, Electronic Journal of Linear Algebra, 20 (2010) 723-729
%\bibitem{Dickinson2011} P.I.C. Dickinson, Geometry of the copositive and completely positive cones, Journal of Mathematical Analysis and Applications, 2011, 380(1): 377-395
%\bibitem{DQW2013} W. Ding, L. Qi and Y. Wei, $M$-tensors and nonsingular $M$-tensors, Linear Algebra and Its Applications 439 (2013) 3264-3278
%\bibitem{DQW2015} W. Ding, L. Qi and Y. Wei, Sum-of-squares decomposition and augmented Vandermonde decomposition of strong Hankel tensors: Theory and algorithms, arXiv preprint arXiv:1505.02528, 2015
%\bibitem{DA2013} H. Dong and K. Anstreicher, Separating doubly nonnegative and completely positive matrices, Mathematical Programming, 2013, 137(1-2): 131-153
%\bibitem{D2010} M. D$\ddot{\text{u}}$r, Copositive programming-a survey, Recent advances in optimization and its applications in engineering. Springer Berlin Heidelberg, 2010: 3-20
%\bibitem{DS2008} M. D$\ddot{\text{u}}$r and G. Still, Interior points of the completely positive cone, Electronic Journal of Linear Algebra, 17 (2008) 48-53
%\bibitem{F2015} J. Fan, CP-decomposition of positive Cauchy tensors, Private Communication, 2015
%\bibitem{FZ2014} J. Fan and A. Zhou, Completely positive tensor decomposition, arXiv preprint arXiv:1411.5149, 2014

%\bibitem{GST2013} M. S. Gowda and R. Sznajder and J. Tao, The automorphism group of a completely positive cone and its Lie algebra, Linear Algebra and its Applications, 2013, 438(10) 3862-3871

%\bibitem{HJ1990} R. A. Horn and C. R. Johnson, Matrix Analysis, Cambridge Univeristy Press, Cambridge, UK, 1990
%\bibitem{HQ2015} S. Hu and L. Qi, The Laplacian of a uniform hypergraph, Journal of Combinatorial Optimization, 29 (2015) 331-366.
%\bibitem{HQS2013} S. Hu, L. Qi and J. Shao, Cored hypergraphs, power hypergraphs and their Laplacian eigenvalues, Linear Algebra and Its Applications 439 (2013) 2980-2998
\bibitem{HQX2015} S. Hu, L. Qi and J. Xie, The largest Laplacian and signless Laplacian eigenvalues of a uniform hypergraph, Linear Algebra and Its Applications, 469 (2015) 1-27
%\bibitem{HS2010} J. B. Hiriart-Urruty and A. Seeger, A variational approach to copositive matrices, SIAM review, 2010, 52(4): 593-629
%\bibitem{KMB2015} M.R. Kannan, N. Shaked-Monderer and A. Berman, Some properties of strong $H$-tensors and general $H$-tensors, Linear Algebra and its Applications, 476 (2015) 42-55
%\bibitem{KKT2013} S. Kim, M. Kojima and K. C. Toh, A Lagrangian-DNN relaxation: a fast method for computing tight lower bounds for a class of quadratic optimization problems, Preprint, 2013 http://www.optimization-online.org/DB\_HTML/2013/10/4073
%\bibitem{K2015} T. Kolda, Numerical optimization for symmetric tensor decomposition, Math. Program., Ser.B, (2015) 151: 225-248
%\bibitem{LDV2000} De Lathauwer L, De Moor B, Vandewalle J. On the best rank-$1$ and rank-$(r_1, r_2,\ldots, r_n)$ approximation of higher-order tensors, SIAM Journal on Matrix Analysis and Applications, 2000, 21(4): 1324-1342
%\bibitem{LDV2001}  L. De Lathauwer, B. De Moor and J. Vandewalle, Independent component analysis and (simultaneous) third-order tensor diagonalization, Signal Processing, IEEE Transactions on, 49(10) (2001) 2262-2271
%\bibitem{LL2015} C. Li and Y. Li, Double $B$-tensors and quasi-double $B$-tensors, Linear Algebra and its Applications 466 (2015) 343-356

%\bibitem{LQL2015} C. Li, L. Qi and Y. Li, $MB$-tensors and $MB_0$ tensors, Preprint, 2015
%\bibitem{LWZZL2014} C. Li, F. Wang, J. Zhao, Y. Zhu and Y. Li, Criterions for the positive definiteness of real supersymmetric tensors, Journal of Computational and Applied Mathematics 255 (2014) 1-14
%\bibitem{LQX2014} G. Li, L. Qi and Y. Xu, SOS-Hankel tensors: Theory and application, arXiv:1410.6989, 2014
%\bibitem{LBAM2004} C. Liu, R. Bammer and B. Acar, Characterizing non-Gaussian diffusion by using generalized diffusion tensors, Magnetic Resonance in Medicine, 2004, 51(5): 924-937
%\bibitem{LQX} Z. Luo, L. Qi and N. Xie, The sparsest solutions to Z-tensor complementarity problems, arXiv:1505.00993, 2015
%\bibitem{LQY2015} Z. Luo, L. Qi and Y. Ye, Linear operators and positive semidefiniteness of symmetric tensor spaces, Science China Mathematics, 58 (2015) 197-212

%\bibitem{PVZ2014} J. Pe$\tilde{\text{n}}$a, J. Vera and L. Zuluaga, Completely positive reformulations for polynomial optimization, Math. Program., Ser. B, (2015) 151: 405-431

\bibitem{Qi2005} L. Qi, Eigenvalues of a real supersymmetric tensor, Journal of Symbolic Computation 40 (2005) 1302-1324
\bibitem{Qi2013} L. Qi, Symmetric nonnegative tensors and copositive tensors, Linear Algebra and Its Applications, 439 (2013) 228-238
\bibitem{Qi2014} L. Qi, $H^+$-eigenvalues of Laplacian and signless Lapaclian tensors, Communications in Mathematical Sciences, 12 (2014) 1045-1064
\bibitem{QSW2014} L. Qi, J. Shao and Q. Wang, Regular uniform hypergraphs,
$s$-cycles, $s$-paths and their largest Laplacian H-eigenvalues, Linear Algebra and Its Applications,
443 (2014) 215-227
\bibitem{QWC2015} L. Qi, Q. Wang and Y. Chen, Three dimensional strongly symmetric circulant tensors, Linear Algebra and Its Applications  482 (2015) 207-220
%\bibitem{Q2015}  L. Qi, Hankel tensors: Associated Hankel matrices and Vandermonde decomposition, Communications in Mathematical Sciences 13 (2015) 113-125
%\bibitem{QS2014} L. Qi and Y. Song, An even order symmetric $B$ tensor is positive definite, Linear Algebra and Its Applications 457 (2014) 303-312
%\bibitem{QSZ} L. Qi, D. Sun and G. Zhou, A new look at smoothing Newton methods for nonlinear complementarity problems and box constrained variational inequalities, Mathematical Programming 87 (2000) 1-35
\bibitem{QXX2014} L. Qi, C. Xu and Y. Xu, Nonnegative tensors factorization, completely positive tensors, and a hierarchical elimination algorithm, SIAM Journal on Matrix Analysis and Applications, 35(4) (2014) 1227-1241
%\bibitem{QY2014} L. Qi and Y. Ye, Space tensor conic programming, Comput Optim Appl, 2014, 59: 307-319
%\bibitem{SQ2014} Yisheng Song and Liqun Qi, Properties of some classes of structured tensors, arXiv:1403.1118v4, 2014
%\bibitem{SZ1993} L. Salce and P. Zanardo, Completely positive matrices and positivity of least squares solutions, Linear algebra and its applications, 178 (1993) 201-216
%\bibitem{SH} A. Shashua and T. Hazan, Non-negative tensor factorization with applications to statistics and computer vision, {\sl ACM International Conference Proceeding Series: Proceedings of the 22nd international conference on Machine learning}, 119  (2005) 792 - 799.
\bibitem{SSW2015} J. Shao, H. Shan, B. Wu, Some spectral properties and characterizations of connected odd-bipartite uniform hypergraphs, to appear in:  Linear and Multilinear Algebra, (2015)
%\bibitem{SongQi2014} Y. Song and L. Qi, Infinite and finite dimensional Hilbert tensors, Linear Algebra Appl.,451 (2014) 1-14
%\bibitem{SQ2014} Y. Song and L. Qi, Properties of tensor complementarity problem and some classes of structured tensors, arXiv:1412.0113v2, 2015
%\bibitem{SQ2015} Y. Song and L. Qi, Tensor complementarity problem and semi-positive tensors, arXiv:1502.02209, 2015

%\bibitem{Xu2015} Changqing Xu, Hankel tensors, Vandermonde tensors and their positivities, Linear Algebra and Its Applications, doi:10.1016/j.laa.2015.02.012
\bibitem{Wa2007} W.D. Wallis, Introduction to Combinatorical Designs, Second Edition, Chapman \& Hall/CRC, New York, 2007.
\bibitem{Wi2015} Wikipedia, Block design, 5 August 2015.
%\bibitem{YM2010} A. Yoshise and Y. Matsukawa, On optimization over the doubly nonnegative cone, 2010 IEEE International Symposium on Computer-Aided Control System Design (CACSD) 13-19
%\bibitem{YY2010} Y. Yang and Q. Yang, Further results for Perron-Frobenius theorem for nonnegative tensors, SIAM Journal on Matrix Analysis and Applications, 31(5) (2010) 2517-2530
%\bibitem{YY2014} Pingzhi Yuan and Lihua You, Some remarks on $P$, $P_0$, $B$ and $B_0$ tensors, Linear Algebra and Its Applications 459 (2014) 511-521

%\bibitem{ZLQ2012} Zhang X, Ling C, Qi L. The best rank-$1$ approximation of a symmetric tensor and related spherical optimization problems, SIAM Journal on Matrix Analysis and Applications, 2012, 33(3): 806-821
%\bibitem{ZQZ2014} L. Zhang, L. Qi and G. Zhou, $M$-tensors and some applications, SIAM Journal on Matrix Analysis and Applications, 35(2) (2014) 437-452
%\bibitem{ZF2014} A. Zhou and J. Fan, The CP-matrix completion problem, SIAM Journal on Matrix Analysis and Applications, 35(1) (2014) 127-142


%\bibitem{PVZ2014} J. Pe$\tilde{n}$a, J, Vera, L. Zuluaga, Completely positive reformulations for polynomial
%optimization, Math. Program., Ser. B, 2014 DOI 10.1007/s10107-014-0822-9

%\bibitem{ZF2014} A. Zhou and J. Fan, The CP-matrix completion problem, SIAM Journal on Matrix Analysis and Applications, 35(1) (2014) 127-142

%\bibitem{K2015} T. Kolda, Numerical Optimization for Symmetric Tensor Decomposition, arXiv preprint arXiv:1410.4536, 2014

%\bibitem{FZ2014} J. Fan and A. Zhou, Completely positive tensor decomposition, arXiv preprint arXiv:1411.5149, 2014

%\bibitem{B2012} I.M. Bomze, Copositive Optimization,In Floudas, C>A., Pardalos, P.M.(eds) Encyclopedia of Oprimization, 2nd edn, pp 561-564, Springer, Berlin (2009)

\end{thebibliography}

\documentclass[twoside,10pt]{article}
\usepackage{amssymb,amsmath,euscript}
\usepackage{graphicx,color,caption}

\usepackage{algorithm,algpseudocode}

\newtheorem{proposition}{Proposition}[section]
\newtheorem{theorem}[proposition]{Theorem}

\newtheorem{definition}[proposition]{Definition}

\newcommand{\qed}{\hphantom{.}\hfill $\Box$\medbreak}

\def\S{\mathbb{S}}
\def\P{\mathcal{P}}
\def\D{\mathcal{D}}

\def\T{\mathbb{T}}
\def\Q{\mathcal{Q}}
\def\O{\mathcal{O}}
\def\A{\mathcal{A}}
\def\J{\mathcal{J}}
\def\B{\mathcal{B}}
\def\C{\mathcal{C}}

\def\I{\mathcal{I}}
\def\J{\mathcal{J}}
\def\x{{\bf x}}
\def\u{{\bf u}}
\def\0{{\bf 0}}
\title{\bf{Characterization Tensors of Balanced Incomplete Block Designs}
\thanks{This research was supported by the National Natural Science Foundation of China (11301022,11431002), the State Key Laboratory of Rail Traffic Control and Safety, Beijing Jiaotong University (RCS2014ZT20, RCS2014ZZ01), and the Hong Kong Research Grant Council (Grant No.
PolyU 502111, 501212, 501913 and 15302114).}}
\author{Liqun Qi \thanks{Department of Applied Mathematics, The Hong Kong Polytechnic University, Hung Hom, Kowloon, Hong Kong. ({\tt
liqun.qi@polyu.edu.hk}).}  \hspace{1mm}   Ziyan
Luo,  \thanks{ State Key Laboratory of Rail Traffic Control and Safety, Beijing
Jiaotong University, Beijing 100044, P.R. China; ({\tt
starkeynature@hotmail.com}).}
}

\begin{document}
\maketitle

\begin{abstract}
Balanced incomplete block designs (BIBDs) have wide applications in engineering, business and sciences.   In this paper, for each $(v, k, \lambda)$-BIBD, we construct  a strongly symmetric $k$-th order $v$-dimensional tensor.    We call such a strongly symmetric tensor the characterization tensor of that BIBD, and the absolute value tensor of
the characterization tensor the signless characterization tensor of that BIBD.   We study some spectral properties of such characterization tensors and signless characterization tensors.  In this way, we provide a new tool to study BIBDs.

\vskip 12pt \noindent {\bf Key words.} {balanced incomplete block designs, characterization tensors, signless characterization tensors, H-eigenvalues}

\vskip 12pt\noindent {\bf AMS subject classifications. }{15A18, 15A69, 15B48}
\end{abstract}


\section{Introduction}
Balanced incomplete block designs (BIBDs) have wide applications in engineering, business and sciences \cite{Wa2007, Wi2015}.

Given a finite set $X$ of points and integers $k, r, \lambda \ge 1$, we may define a {\bf balanced incomplete block design} $B$ to be a family of $k$-element subsets of $X$, called {\bf blocks} such that for each $i \in X$, there are exactly $r$ blocks containing $i$, and for each pair $(i, j)$ of points in $X$, there are exactly $\lambda$ blocks containing $i$ and $j$.    Let the cardinality of $X$ to be $v$ and the cardinality of $B$ be $b$.   Then we have
$$bk = vr$$
and
$$\lambda(v-1) = r(k-1).$$
We assume the design is simple, i.e., repeated blocks are not allowed.   We also assume that $v > k$.  In design, this is the meaning of the word ``incomplete'' \cite{Wa2007, Wi2015}.

In this paper, for each $(v, k, \lambda)$-BIBD, we construct a strongly symmetric $k$-th order $v$-dimensional tensor.    We call such a strongly symmetric tensor the characterization tensor of that BIBD, and the absolute value tensor of
the characterization tensor the signless characterization tensor of that BIBD.   We study some spectral properties of such characterization tensors and signless characterization tensors.  In this way, we provide a new tool to study BIBDs.

Some notations that will be used throughout the paper are listed here. The $v$-dimensional real Euclidean space is denoted by $\Re^v$, where $v$ is a given natural number. The nonnegative orthant in $\Re^v$ is denoted by $\Re^v_+$, with the interior $\Re^v_{++}$ consisting of all positive vectors. 
Denote $[v]:=\{1,2,\ldots, v\}$. Vectors are denoted by bold letters such as $\x$, $\u$, matrices are denoted by capital letters such as $A$, $P$, and tensors are written as calligraphic capital letters such as $\A$, $\B$. The space of all real $k$th order $v$-dimensional tensors is denoted by $\T_{k,v}$, and the space of all symmetric tensors in $\T_{k,v}$ is denoted by $\S_{k,v}$. 
For a subset $\Gamma\subseteq [v]$, $|\Gamma|$ stands for its cardinality.



\section{Strongly Symmetric Tensors and Eigenvalues}
Let $\A=\left(a_{i_1\ldots i_k}\right)$ be a $k$th order $v$-dimensional real tensor. $\A$ is called a \emph{symmetric} tensor if the entries $a_{i_1\ldots i_k}$ are invariant under any permutation of their indices for all $i_j\in [v]$ and $j\in [k]$, denoted as $\A\in \S_{k,v}$. A symmetric tensor $\A$ is said to be positive semidefinite (definite) if $\A \x^k:=\sum\limits_{i_1,\ldots,i_k\in [v]}a_{i_1\ldots i_k}x_{i_1}\cdots x_{i_k}\geq 0 (>0)$ for any $\x\in \Re^v\setminus \{\0\}$ \cite{Qi2005}. Here, $\x^k$ is a rank-one tensor in $\S_{k,v}$ defined as $\left(\x^k\right)_{i_1\ldots i_k}:=x_{i_1}\cdots x_{i_k}$ for all $i_1$, $\ldots$, $i_k\in [v]$. Evidently, when $k$ is odd, $\A$ could not be positive definite and $\A$ is positive semidefinite if and only if $\A=\O$, where $\O$ stands for the zero tensor. A tensor $\A\in \S_{k,v}$ is said to be (strictly) copositive if $\A x^k\geq 0$ ($>0$) for all $x\in\Re^n_{+}\setminus \{\0\}$ \cite{Qi2013}. The definitions on eigenvalues of symmetric tensors are as follows.

\begin{definition}[\cite{Qi2005}] Let $\A\in \S_{k,v}$ and $\mathbb{C}$ be the complex field. We say that $(\mu, \x)\in {\mathbb{C}}\times \left({\mathbb{C}}^v\setminus \{\0\}\right)$ is an eigenvalue-eigenvector pair of $\A$ if
$\A \x^{k-1}=\mu x^{[k-1]},$ where $\A \x^{k-1}$ and $\x^{[k-1]}$ are all $v$-dimensional column vectors given by
$$\left(\A \x^{k-1}\right)_i:=\sum\limits_{i_2,\ldots,i_k\in[v]} a_{ii_2\ldots i_k} x_{i_2}\cdots x_{i_k},~\left(\x^{[k-1]}\right)_i=x_i^{k-1},~~\forall i\in [v].\eqno(2.1)$$  If the eigenvalue $\mu$ and the eigenvector $\x$ are real, then $\mu$ is called an $H$-eigenvalue of $\A$ and $\x$ an $H$-eigenvector of $\A$ associated with $\mu$. If $\x\in \Re^n_+ (\Re^n_{++})$, then $\mu$ is called an $H^+ (H^{++})$-eigenvalue of $\A$.  The maximum modulus of the eigenvalues of $\A$ is called the spectral radius of $\A$ and denoted by $\rho(\A)$.   The largest $H$-eigenvalue of $\A$ is denoted as $\mu_{\max}(\A)$.  The set of all the $H$-eigenvalues of $\A$ is called the $H$-spectrum of $\A$.
\end{definition}

When $k$ is even, $\A\in \S_{k,v}$ always has $H$-eigenvalues \cite{Qi2005}.   A nonnegative tensor always has $H$-eigenvalues \cite{CQZ2013}.

\begin{definition}[\cite{Qi2005}] Let $\A\in \S_{k,v}$ and $\mathbb{C}$ be the complex field. We say that $(\mu,\x)\in {\mathbb{C}}\times \left({\mathbb{C}}^v\setminus \{\0\}\right)$ is an $E$ eigenvalue-eigenvector pair of $\A$ if
$\A \x^{k-1}=\mu \x$ and $\x^{T}\x=1$, where $\A \x^{k-1}$ is defined as in (2.1). If the $E$-eigenvalue $\mu$ and the eigenvector $\x$ are real, then $\mu$ is called a $Z$-eigenvalue of $\A$ and $\x$ a $Z$-eigenvector of $\A$ associated with $\mu$.
\end{definition}

$\A\in \S_{k,v}$ always has $Z$-eigenvalues \cite{Qi2005}.

A $k$th order $v$-dimensional real tensor $\A=\left(a_{i_1\ldots i_k}\right)$ is called a \emph{strongly symmetric} tensor if $a_{i_1\ldots i_k} = a_{j_1\ldots j_k}$ as long as
$\{ i_1, \ldots, i_k \} = \{ j_1, \ldots, j_k \}$ for all $i_l, j_l\in [v]$ and $l\in [k]$.   For example, for a fourth order $v$-dimensional real tensor $\A=\left(a_{i_1i_2i_3i_4}\right)$, we always have $a_{ijjj}=a_{ijij}=a_{jiii}$ for all $i, j \in [v]$.  Denote the set of all $k$th order $v$-dimensional real strongly symmetric tensors as $\S\S_{k,v}$.  Strongly symmetric tensors were introduced in \cite{QXX2014} and further studied in \cite{QWC2015}.

Let $\A=\left(a_{i_1\ldots i_k}\right) \in \T_{k,v}$.    We call $a_{i_1\ldots i_k}$ a diagonal entry of $\A$ if $i_1= \ldots = i_k$.   Otherwise, we call it an off-diagonal entry of $\A$.
If $| \{ i_1, \ldots, i_k \} | = 2$, we say that $a_{i_1\ldots i_k}$ is a sub-diagonal entry of $\A$.   A tensor in $\T_{k,v}$ is called a diagonal tensor if all of its off-diagonal entries are $0$.  A diagonal tensor in $\T_{k,v}$ is called an identity tensor of $\T_{k,v}$, and denoted by $\I$ if all of its diagonal entries are $1$.  A tensor in $\T_{k,v}$ is called a sub-diagonal tensor if all of its entries, which are not sub-diagonal entries are $0$.   A sub-diagonal tensor in $\T_{k,v}$ is called an identity sub-diagonal tensor of $\T_{k,v}$, and denoted by $\J$ if all of its sub-diagonal entries are $1$ and the other entries are $0$.  Clearly, diagonal tensors and sub-diagonal tensors are strongly symmetric tensors.

\section{Characterization Tensors of Balanced Incomplete Block Designs}

We may regard $(v, k, \lambda)$-BIBD as a $k$-uniform hypergraph $G = (X, B)$ \cite{CD2012, Qi2014}, where $X$ is the vertex set and $B$ is the edge set.  For hypergraphs, for each vertex $i \in X$, let $X_i = \{ e \in B : i \in e \}$.  Then $d(i) = |X_i|$ is called the degree of $i$.   For each pair of vertices $i, j \in X$, let $X_{i, j} = \{ e \in B : i, j \in e \}$.  Then $d(i, j) = |X_{i, j}|$ is called the co-degree of $(i, j)$.   A $k$-uniform hypergraph is called a $r$-regular $k$-uniform hypergraph if the degrees of its vertices are the same as $r$ \cite{CD2012, QSW2014}.  Thus, a $(v, k, \lambda)$-BIBD is corresponding to a $r$-regular $k$-uniform hypergraph with the co-degrees of its vertices are the same as $\lambda$.

For a $k$-uniform hypergraph $G = (X, B)$, there are several tensors associated with it.  The adjacency tensor $\A$ of $G$ is defined as a tensor $\A=\left(a_{i_1\ldots i_k}\right) \in \S_{k,v}$, with $a_{i_1\ldots i_k} = {1 \over (k-1)!}$ if $(i_1, \ldots, i_k) \in B$, and $0$ otherwise \cite{CD2012}.   The degree tensor $\D$ of $G$ is defined as a diagonal tensor in $\S_{k,v}$ with its $i$th diagonal entries as $d(i)$.   The Laplacian tensor and signless Laplacian tensor of $G$ are defined as $\D - \A$ and $\D + \A$ respectively \cite{Qi2014}.  Spectral hypergraph theory via tensors deal with the spectral properties of adjacency tensors, Laplacian tensors and signless Laplacian tensors, and their relations with hypergraphs \cite{CCL2015, CD2012, HQX2015, Qi2014, QSW2014, SSW2015}.  We now define the co-degree tensor $\C=(c_{i_1\ldots i_k})$ of $G$ by $c_{i_1\ldots i_k}= {d(i, j) \over 2^{k-1}-1}$ if $\{ i_1, \ldots, i_k \} = \{ i, j \}, i \not = j$, and $0$ otherwise.  Thus, $G$ is corresponding to a $(v, k, \lambda)$-BIBD if its degree tensor $\D = r\I$ and its co-degree tensor $\C = {\lambda \over 2^{k-1}-1}\J$.
Clearly, all these tensors involved are strongly symmetric tensors.

Now, for a $(v, k, \lambda)$-BIBD, we may define its {\bf characterization tensor} as
$$\P = 2r(k-1)\I + \C - (k-1)\A$$
and its {\bf signless characterization tensor} as
$$\Q = 2r(k-1)\I + \C  + (k-1)\A.$$

Since $\lambda \ge 1$, the hypergraph $G$ corresponding to a $(v, k, \lambda)$-BIBD is connected.  Recall that \cite{HQX2015} a $k$-uniform hypergraph $G = (X, B)$ is called odd-bipartite if $k$ is even and there is a subset $Y$ of $X$ such that for every $e \in B$, $| e \cap Y|$ is odd.   A $(v, k, \lambda)$-BIBD is called odd-bipartite if $k$ is even and the hypergraph $G$ corresponding to it is odd-bipartite.

We now have the following theorem on the spectral properties of $\P$ and $\Q$ of a $(v, k, \lambda)$-BIBD.

\medskip

\begin{theorem}
Suppose that $k, r, \lambda \ge 1$ and $v \ge k$.   Let $\P$ and $\Q$ be the characterization and signless characterization tensors of a $(v, k, \lambda)$-BIBD.   Then

(a) Any eigenvalue $\mu$ of $\P$ and $\Q$ satisfies
$$| \mu -  2r(k-1) | \le 2r(k-1).$$

(b) $\rho(\Q) = 4r(k-1)$.

(c) $\P$ and $\Q$ are co-positive.   When $k$ is even, they are positive semi-definite.

(d) $\mu_{\max}(\P) \le mu_{\max}(\Q)$.  The equality holds here if and only if $k$ is even and the BIBD is odd-bipartite.   If $k$ is even and the BIBD is odd-bipartite, then $\P$ has $H$-eigenvalues and the $H$-spectrum of $\P$ is equal to the $H$-spectrum of $\Q$.
\end{theorem}

{\bf Proof. }  By the Gershgorin theorem for tensor eigenvalues and the definitions of $\P$ and $\Q$, we have (a).

Let $\u$ be the all $1$ vector in $\Re^v$.  Then we have $\Q \u^{k-1} = 4r(k-1)\u^{[k-1]}$.   By (a) and the theory of nonnegative tensors \cite{CQZ2013}, we have (b).

By (a) and \cite{Qi2013, Qi2005}, we have (c).

With a way similar to \cite{HQX2015, SSW2015}, we may prove (d).
\qed

The question is, can we derive some performance properties from the spectral properties of $\P$ and $\Q$.

\section*{Acknowledgements}
The first author is thankful to Prof. Sanming Zhou, who introduced the topic of block design to him, and to Prof. Linyuan Lu for the discussion on co-degrees.


\end{document}